# Commonality in Few: Few-Shot Multimodal Anomaly Detection via Hypergraph-Enhanced Memory

### v1

@@ Introduction
We denote the test hypergraph as $\mathcal{G}^{test}=(X^{test}\in\mathbb{R}^{N\times dim},\mathbf{H}^{test}\in\{0,1\}^{N\times\mathcal{E}_{num}})$ and the memory bank hypergraph as $\mathcal{G}^{mem}=(X^{mem}\in\mathbb{R}^{A\times dim},\mathbf{H}^{mem}\in\{0,1\}^{A\times\mathcal{E}_{num}})$, where dimdim is the feature dimension of each node, $\mathbf{H}^{*}$ is the incidence matrix, $N$ is the number of nodes in the test sample, and $A$ is the number of nodes in the memory bank.
We compute the cosine similarity between $X^{test}$ and $X^{mem}$, and for each test node $x^{test}_{i}$, we select the top-$k$ most similar memory bank nodes. We construct a hyperedge that includes $x^{test}_{i}$ and the selected $k$ memory bank nodes. In total, $N$ such cross-domain hyperedges are constructed, resulting in an incidence matrix $\mathbf{H}^{t\rightarrow m}\in\{0,1\}^{(A+N)\times N}$. Similarly, for each memory node $x^{mem}_{i}$, we select the top-$k$ most similar test nodes, and construct a hyperedge that includes $x^{mem}_{i}$ and the selected $k$ test nodes. In total, $A$ such cross-domain hyperedges are constructed, resulting in an incidence matrix $\mathbf{H}^{m\rightarrow t}\in\{0,1\}^{(A+N)\times A}$. By concatenating $\mathbf{H}^{t\rightarrow m}$ and $\mathbf{H}^{m\rightarrow t}$ along the hyperedge dimension, we obtain the cross-domain incidence matrix $\mathbf{H}^{cross}\in\{0,1\}^{(A+N)\times(A+N)}$.
@@ Acknowledgments
To mitigate background interference during detection, we follow AST (rudolph2023asymmetric) and use 3D point clouds to extract foreground masks. We first obtain the depth channel of the point cloud, and fill missing depth values with $3$ iterations of neighborhood mean interpolation after normalization. The background is estimated from the corner depth values, the distances between all other areas and this background are computed, and extract the foreground mask by setting the threshold to $7\times 10^{-3}$.
@@
Table 5 presents the anomaly detection (I-AUROC) and anomaly localization (AUPRO) results for all categories in MVTec 3D-AD in Section 4.4.

### v2

@@ Introduction
We denote the test hypergraph as $\mathcal{G}^{test}=(X^{test}\in\mathbb{R}^{N\times dim},\mathbf{H}^{test}\in\{0,1\}^{N\times\mathcal{E}_{num}})$ and the memory bank hypergraph as $\mathcal{G}^{mem}=(X^{mem}\in\mathbb{R}^{A\times dim},\mathbf{H}^{mem}\in\{0,1\}^{A\times\mathcal{E}_{num}})$, where dimdim is the feature dimension of each node, $\mathbf{H}^{*}$ is the incidence matrix, NN is the number of nodes in the test sample, and AA is the number of nodes in the memory bank.
We compute the cosine similarity between $X^{test}$ and $X^{mem}$, and for each test node $x^{test}_{i}$, we select the top-$k$ most similar memory bank nodes. We construct a hyperedge that includes $x^{test}_{i}$ and the selected $k$ memory bank nodes. In total, NN such cross-domain hyperedges are constructed, resulting in an incidence matrix $\mathbf{H}^{t\rightarrow m}\in\{0,1\}^{(A+N)\times N}$. Similarly, for each memory node $x^{mem}_{i}$, we select the top-$k$ most similar test nodes, and construct a hyperedge that includes $x^{mem}_{i}$ and the selected $k$ test nodes. In total, AA such cross-domain hyperedges are constructed, resulting in an incidence matrix $\mathbf{H}^{m\rightarrow t}\in\{0,1\}^{(A+N)\times A}$. By concatenating $\mathbf{H}^{t\rightarrow m}$ and $\mathbf{H}^{m\rightarrow t}$ along the hyperedge dimension, we obtain the cross-domain incidence matrix $\mathbf{H}^{cross}\in\{0,1\}^{(A+N)\times(A+N)}$.
@@ Acknowledgments
To mitigate background interference during detection, we follow AST (rudolph2023asymmetric) and use 3D point clouds to extract foreground masks. We first obtain the depth channel of the point cloud, and fill missing depth values with 33 iterations of neighborhood mean interpolation after normalization. The background is estimated from the corner depth values, the distances between all other areas and this background are computed, and extract the foreground mask by setting the threshold to $7\times 10^{-3}$.
@@
Table 5 presents the anomaly detection (I-AUROC) and anomaly localization (AUPRO) results for all categories in MVTec 3D-AD in Section $4.$